# Bandits with Preference Feedback: A Stackelberg Game Perspective

### v1

@@ 2 Related Work
Referred to as dueling bandits, a rich body of work considers (finite) multi-armed domains and learns a preference matrix specifying the relation among the arms. Such work often relies on efficient sorting or tournament systems based on the frequency of wins for each action (Jamieson and Nowak, 2011; Zoghi et al., 2014b; Falahatgar et al., 2017). Rather than jointly selecting the arms, such strategies often simplify the problem by selecting one at random (Zoghi et al., 2014a; Zimmert and Seldin, 2018), greedily (Chen and Frazier, 2017), or from the set of previously selected arms (Ailon et al., 2014). In contrast, we jointly optimize both actions by choosing them as the equilibrium of a two-player zero-sum Stackelberg game, enabling a more efficient exploration/exploitation trade-off.
@@ 3 Problem Setting
Consider an agent which repeatedly interacts with an environment: at step $t$ the agent selects two actions ${\bm{x}}_{t},\,{\bm{x}}^{\prime}_{t}\in{\mathcal{X}}$ and only observes stochastic binary feedback $y_{t}\in[0,1]$ indicating if ${\bm{x}}_{t}\succ{\bm{x}}_{t}^{\prime}$, that is, if action ${\bm{x}}_{t}$ is preferred over action ${\bm{x}}^{\prime}_{t}$. Formally, ${\mathbb{P}}(y_{t}=1|{\bm{x}}_{t},{\bm{x}}^{\prime}_{t})={\mathbb{P}}({\bm{x}}_{t}\succ{\bm{x}}_{t}^{\prime})$, and $y_{t}=0$ with probability $1-{\mathbb{P}}({\bm{x}}_{t}\succ{\bm{x}}_{t}^{\prime})$. Based on the preference history $H_{t}=\{({\bm{x}}_{1},{\bm{x}}^{\prime}_{1},y_{1}),\dots({\bm{x}}_{t},{\bm{x}}^{\prime}_{t},y_{t})\}$, the agent aims to sequentially select favorable action pairs. Over a horizon of $T$ steps, the success of the agent is measured through the cumulative dueling regret
@@
where $s:{\mathbb{R}}\rightarrow[0,1]$ is the sigmoid function, i.e. $s(a)=(1+e^{-a})^{-1}$. Referred to as the Bradley-Terry model (Bradley and Terry, 1952), this probabilistic model for binary feedback is widely used in the literature for logistic and generalized bandits (Filippi et al., 2010; Faury et al., 2020). Under the utility-based model, ${\bm{x}}^{\star}=\operatorname*{arg\,max}_{{\bm{x}}\in{\mathcal{X}}}f({\bm{x}})$ and we can draw connections to a classic bandit problem with direct feedback over the utility $f$. In particular, Saha (2021) shows that the dueling regret of (1) is equivalent up to constant factors, to the average utility regret of the two actions, that is $\sum_{t=1}^{T}{f}({\bm{x}}^{\star})-[{f}({\bm{x}}_{t})+{f}({\bm{x}}_{t}^{\prime})]/2$.
@@ 5.2 Action Selection Strategy
We propose MaxMinLCB in Algorithm 1 for the preference feedback bandit problem that selects ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ jointly as
@@ 6 Experiments
Our experiments are on finding the maxima of test functions commonly used in (non-convex) optimization literature (Jamil and Yang, 2013), given only preference feedback. These functions cover challenging optimization landscapes including several local optima, plateaus, and valleys, allowing us to test the versatility of MaxMinLCB. We use the Ackley function for illustration in the main text and provide the regret plots for the remainder of the functions in Appendix E. For all experiments, we set the horizon $T=2000$ and evaluate all algorithms on a uniform mesh over the input domain of size $100$. All experiments are run across $20$ random seeds and reported values are averaged over the seeds, together with standard error. The environments and algorithms are implemented [^1] The code is made available at github.com/lasgroup/MaxMinLCB. end-to-end in JAX (Bradbury et al., 2018).
@@ 6.2 Benchmarking Acquisition Functions
In this section, we compare MaxMinLCB with other utility-based bandit algorithms. To isolate the benefits of our acquisition function, we instantiate all algorithms with the same confidence sets, and use our improved preferred-based bound of Corollary 5. Therefore, our implementation differs from the corresponding references, while we refer to the algorithms by their original name. We consider the following baselines. Doubler and MultiSBM (Ailon et al., 2014) choose ${\bm{x}}_{t}$ as a reference action from the recent history of actions and pair it with ${\bm{x}}^{\prime}_{t}$ which maximizes the joint UCB (cf. Algorithm 5 and 6). RUCB (Zoghi et al., 2014a) chooses ${\bm{x}}^{\prime}_{t}$ similarly, however, it selects the reference action uniformly at random from ${\mathcal{M}}_{t}$ (Algorithm 7). MaxInP (Saha, 2021) also maintains the set of plausible maximizers ${\mathcal{M}}_{t}$, and at each time step, it selects the pair of actions that maximize $\sigma_{t}^{D}({\bm{x}},{\bm{x}}^{\prime})$ (Algorithm 3). IDS (Kirschner and Krause, 2021) selects the reference action greedily by maximizing $f_{t}$, and pairs it with an informative action (Algorithm 4). Notably, all algorithms, with the exception of MaxInP, choose one of the actions independently and use it as a reference point when selecting the other one. Figure 2 illustrates the differences in action selection between UCB, maximum information, and MaxMinLCB approaches.
@@
Table 1: Benchmarking $R^{\mathrm{D}}_{T}$ for a variety of test utility functions, $T=2000$.
$f$ MaxMinLCB Doubler MultiSBM MaxInP RUCB IDS
Ackley $\boldsymbol{54}\pm 3$ $67\pm 3$ $453\pm 58$ $112\pm 5$ $79\pm 3$ $99\pm 10$
Branin $\boldsymbol{63}\pm 10$ $79\pm 8$ $213\pm 28$ $197\pm 23$ $\boldsymbol{63}\pm 11$ $86\pm 17$
Eggholder $\boldsymbol{100}\pm 7$ $132\pm 8$ $435\pm 56$ $179\pm 21$ $155\pm 24$ $123\pm 13$
Hoelder $\boldsymbol{107}\pm 16$ $132\pm 8$ $460\pm 59$ $169\pm 15$ $153\pm 18$ $119\pm 15$
Matyas $81\pm 8$ $87\pm 8$ $209\pm 27$ $100\pm 8$ $79\pm 7$ $\boldsymbol{58}\pm 8$
Michalewicz $\boldsymbol{108}\pm 10$ $149\pm 11$ $473\pm 61$ $196\pm 25$ $184\pm 28$ $154\pm 19$
Rosenbrock $\boldsymbol{18}\pm 3$ $24\pm 8$ $131\pm 17$ $76\pm 6$ $38\pm 6$ $34\pm 9$
@@ Appendix A Proofs for Bandits with Logistic Feedback
To analyze our function-valued confidence sequences, we start by re-writing the logistic loss function as
@@ Appendix E Additional Experiments
In this section, we provide Table 2 that details our logistic bandit benchmark, complementing the results in Section 6.1. Figure 3 to Figure 8 show the logistic and dueling regret of additional test functions, complementing the results of Table 1.
@@
Figure 4: Regret with Eggholder utility function with logistic and preference feedback.
Figure 5: Regret with Hölder utility function with logistic and preference feedback.
Figure 6: Regret with Matyas utility function with logistic and preference feedback.
Figure 7: Regret with Michalewicz utility function with logistic and preference feedback.
Figure 8: Regret with Rosenbrock utility function with logistic and preference feedback.

### v2

@@ 2 Related Work
Referred to as dueling bandits, a rich body of work considers (finite) multi-armed domains and learns a preference matrix specifying the relation among the arms. Such work often relies on efficient sorting or tournament systems based on the frequency of wins for each action (Jamieson and Nowak, 2011; Zoghi et al., 2014b; Komiyama et al., 2015; Wu and Liu, 2016; Falahatgar et al., 2017). Rather than jointly selecting the arms, such strategies often simplify the problem by selecting one at random (Zoghi et al., 2014a; Zimmert and Seldin, 2018), greedily (Chen and Frazier, 2017), or from the set of previously selected arms (Ailon et al., 2014). In contrast, we jointly optimize both actions by choosing them as the equilibrium of a two-player zero-sum Stackelberg game, enabling a more efficient exploration/exploitation trade-off.
@@ 3 Problem Setting
Consider an agent which repeatedly interacts with an environment: at step $t$ the agent selects two actions ${\bm{x}}_{t},\,{\bm{x}}^{\prime}_{t}\in{\mathcal{X}}$ and only observes stochastic binary feedback $y_{t}\in\{0,1\}$ indicating if ${\bm{x}}_{t}\succ{\bm{x}}_{t}^{\prime}$, that is, if action ${\bm{x}}_{t}$ is preferred over action ${\bm{x}}^{\prime}_{t}$. Formally, ${\mathbb{P}}(y_{t}=1|{\bm{x}}_{t},{\bm{x}}^{\prime}_{t})={\mathbb{P}}({\bm{x}}_{t}\succ{\bm{x}}_{t}^{\prime})$, and $y_{t}=0$ with probability $1-{\mathbb{P}}({\bm{x}}_{t}\succ{\bm{x}}_{t}^{\prime})$. Based on the preference history $H_{t}=\{({\bm{x}}_{1},{\bm{x}}^{\prime}_{1},y_{1}),\dots({\bm{x}}_{t},{\bm{x}}^{\prime}_{t},y_{t})\}$, the agent aims to sequentially select favorable action pairs. Over a horizon of $T$ steps, the success of the agent is measured through the cumulative dueling regret
@@
where $s:{\mathbb{R}}\rightarrow[0,1]$ is the sigmoid function [^1] May be generalized to differentiable monotonically increasing functions satisfying $s(x)+s(-x)=1$., i.e. $s(a)=(1+e^{-a})^{-1}$. Referred to as the Bradley-Terry model (Bradley and Terry, 1952), this probabilistic model for binary feedback is widely used in the literature for logistic and generalized bandits (Filippi et al., 2010; Faury et al., 2020). Under the utility-based model, ${\bm{x}}^{\star}=\operatorname*{arg\,max}_{{\bm{x}}\in{\mathcal{X}}}f({\bm{x}})$ and we can draw connections to a classic bandit problem with direct feedback over the utility $f$. In particular, Saha (2021) shows that the dueling regret of (1) is equivalent up to constant factors, to the average utility regret of the two actions, that is $\sum_{t=1}^{T}{f}({\bm{x}}^{\star})-[{f}({\bm{x}}_{t})+{f}({\bm{x}}_{t}^{\prime})]/2$.
@@
While our theoretical framework targets euclidean domains, our methodology may be used on general domains of text or images, given vector embeddings obtained via unsupervised learning. Solving a bandit problem on top of embeddings from a pretrained language model is common practice in further fine-tuning of such models (e.g., Nguyen et al., 2024; Mehta et al., 2023a), and is further demonstrated in our Yelp experiment (c.f Section 6.3). Lastly, we highlight that our results may be smoothly extended to contextual bandits with stochastic context, by simply modifying the signature of the kernel function to $k^{\prime}({\bm{x}},{\bm{x}}^{\prime},{\bm{z}}):{\mathcal{X}}\times{\mathcal{X}}\times{\mathcal{Z}}\rightarrow{\mathbb{R}}$, where ${\bm{z}}\in{\mathcal{Z}}$ denotes the context. This setting accommodates applications in active learning for fine-tuning of large language models, where the context is the prompt and the two actions are two alternative responses.
@@ 5.2 Action Selection Strategy
Let ${\bm{x}}^{\prime}({\bm{x}})=\arg\min_{{\bm{x}}^{\prime}\in{\mathcal{M}}_{t}}\mathrm{LCB}_{t}({\bm{x}},{\bm{x}}^{\prime})$ denote a response function. We propose MaxMinLCB in Algorithm 1 for the preference feedback bandit problem that selects ${\bm{x}}_{t}$ and ${\bm{x}}_{t}^{\prime}$ jointly as
@@ 6 Experiments
Our experiments are on finding the maxima of test functions commonly used in (non-convex) optimization literature (Jamil and Yang, 2013), given only preference feedback. These functions cover challenging optimization landscapes including several local optima, plateaus, and valleys, allowing us to test the versatility of MaxMinLCB. We use the Ackley function for illustration in the main text and provide the regret plots for the remainder of the functions in Appendix E. For all experiments, we set the horizon $T=2000$ and evaluate all algorithms on a uniform mesh over the input domain of size $100$. Additionally, we conducted experiments on the Yelp restaurant review dataset to demonstrate the applicability of MaxMinLCB on real-world data and its scaling to larger domains. All experiments are run across $20$ random seeds and reported values are averaged over the seeds, together with standard error. The environments and algorithms are implemented [^2] The code is made available at github.com/lasgroup/MaxMinLCB. end-to-end in JAX (Bradbury et al., 2018).
@@ 6.2 Benchmarking Acquisition Functions
Table 1: Benchmarking $R^{\mathrm{D}}_{T}$ for a variety of test utility functions, $T=2000$. The top 3 rows show results for smoother functions without steep gradients and local optima while the bottom 5 rows show the results for more challenging problems.
$f$ MaxMinLCB Doubler MultiSBM MaxInP RUCB IDS
Branin $\boldsymbol{104}\pm 13$ $114\pm 9$ $\boldsymbol{89}\pm 13$ $340\pm 2$ $\boldsymbol{101}\pm 14$ $163\pm 22$
Matyas $125\pm 5$ $136\pm 4$ $\boldsymbol{106}\pm 7$ $136\pm 6$ $\boldsymbol{106}\pm 6$ $128\pm 5$
Rosenbrock $\boldsymbol{27}\pm 4$ $44\pm 12$ $\boldsymbol{25}\pm 5$ $109\pm 2$ $58\pm 7$ $58\pm 13$
Ackley $\boldsymbol{56}\pm 2$ $72\pm 2$ $65\pm 0.5$ $120\pm 1$ $84\pm 0.7$ $111\pm 9$
Eggholder $\boldsymbol{113}\pm 6$ $154\pm 4$ $134\pm 3$ $230\pm 34$ $213\pm 40$ $141\pm 12$
Hoelder $\boldsymbol{141}\pm 26$ $\boldsymbol{154}\pm 3$ $\boldsymbol{136}\pm 15$ $204\pm 20$ $200\pm 28$ $\boldsymbol{132}\pm 15$
Michalewicz $\boldsymbol{138}\pm 14$ $183\pm 11$ $\boldsymbol{155}\pm 10$ $260\pm 40$ $269\pm 46$ $188\pm 21$
Yelp $\boldsymbol{175}\pm 22$ $263\pm 28$ $\boldsymbol{199}\pm 25$ $409\pm 15$ $214\pm 22$ $255\pm 22$
In this section, we compare MaxMinLCB with other utility-based bandit algorithms. To isolate the benefits of our acquisition function, we instantiate all algorithms with the same confidence sets, and use our improved preferred-based bound of Corollary 5. Therefore, our implementation differs from the corresponding references, while we refer to the algorithms by their original name. We consider the following baselines. Doubler and MultiSBM (Ailon et al., 2014) choose ${\bm{x}}_{t}$ as a reference action from the recent history of actions and pair it with ${\bm{x}}^{\prime}_{t}$ which maximizes the joint UCB (cf. Algorithm 5 and 6). RUCB (Zoghi et al., 2014a) chooses ${\bm{x}}^{\prime}_{t}$ similarly, however, it selects the reference action uniformly at random from ${\mathcal{M}}_{t}$ (Algorithm 7). MaxInP (Saha, 2021) also maintains the set of plausible maximizers ${\mathcal{M}}_{t}$, and at each time step, it selects the pair of actions that maximize $\sigma_{t}^{D}({\bm{x}},{\bm{x}}^{\prime})$ (Algorithm 3). IDS (Kirschner and Krause, 2021) selects the reference action greedily by maximizing $f_{t}$, and pairs it with an informative action (Algorithm 4). Notably, all algorithms, with the exception of MaxInP, choose one of the actions independently and use it as a reference point when selecting the other one. Figure 3 illustrates the differences in action selection between UCB, maximum information, and MaxMinLCB approaches. We note that POP-BO (Xu et al., 2024) and MultiSBM only differ in the estimation of the confidence set. Since we deploy the same confidence set for all acquisition functions, the two algorithms are equivalent and we use MultiSBM in our results, however, comparisons hold for POP-BO as well.
@@
6.3 Real-world Experiment
To further demonstrate the scalability and applicability of MaxMinLCB, we conduct an experiment on the Yelp open dataset of restaurant reviews that we use to design a complex real-world problem. The goal of this experiment is to identify restaurants for each user that fit their respective preferences the best via sequential recommendations. The final dataset after cleaning consists of $275$ restaurants and $20$ users. We define the action space ${\mathcal{X}}$ by assigning to each restaurant their respective $32$-dimensional embedding of their reviews, i.e., ${\mathcal{X}}\subseteq{\mathbb{R}}^{32}$. The dataset provides utility values for users in the form of ratings on the scale of $1$ to $5$, however, not all users rated every restaurant. We estimate missing ratings using collaborative filtering (Schafer et al., 2007). Further details on the data processing are deferred to Section D.1.
Figure 2 shows that the results of this larger problem align with previous conclusions. MaxMinLCB remains the best-performing algorithm with MultiSBM following second. The cumulative regret is also reported in Table 1. Note that neither of the algorithms is tuned or modified for this experiment. These results are only intended to demonstrate that 1) the computations easily scale, and 2) the kernelized approach is still applicable in a text-based domain, by using high-quality vector embeddings.
Figure 2: Regret of learning restaurant recommendations using the Yelp open dataset with preference feedback. Algorithms with different acquisition functions, all using our confidence sets. MaxMinLCB is more sample-efficient.
@@ Appendix A Proofs for Bandits with Logistic Feedback
To analyze our function-valued confidence sequences, we start by re-writing the logistic loss function
@@ Appendix D Details of Experiments
D.1 Yelp Experiment
We filter the Yelp data [^3] Source: https://www.yelp.com/dataset for restaurants in Philadelphia, USA with at least $500$ reviews and users who reviewed at least $90$ restaurants. The final dataset includes $275$ restaurants, $20$ users, and a total of $2563$ reviews. We define the action space ${\mathcal{X}}$ by assigning to each restaurant their respective $32$-dimensional embedding of their reviews, i.e., ${\mathcal{X}}\subseteq{\mathbb{R}}^{32}$. For each restaurant, we concatenate all reviews in the filtered dataset and we use the text-embedding-3-large OpenAI embedding model  [^4] Source: https://platform.openai.com/docs/guides/embeddings to retrieve the embeddings. The Yelp dataset provides utility values for users in the form of ratings on the scale of $1$ to $5$, however, not all users rated every restaurant. We use collaborative filtering to estimate the missing reviews [Schafer et al., 2007]. For each user, these values are used as the utilities for restaurants during the simulation. Experiments are conducted separately for each user, therefore, utility values are not aggregated but the action space is identical for each experiment.
Note that we do not assume any explicit functional form of the utility functions $f$ that we calibrate to this data. Instead, the actions space ${\mathcal{X}}$ and utility values are derived separately from the dataset. Regardless, the results presented in Section 6.3 show that our kernelized method achieves good performance on this task.
@@ Appendix E Additional Experiments
In this section, we provide Table 2 that details our logistic bandit benchmark, complementing the results in Section 6.1. Figure 4 and Figure 9 show the logistic and dueling regret of additional test functions, complementing the results of Table 1.
@@
Figure 9: Top to bottom Regret for Eggholder, Hölder, Matyas Michalewicz, Rosenbrock functions, with logistic and preference feedback.